OpenStudy (gabylovesyou):

The formula for degrees Celsius is: C = 5/9 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (gabylovesyou):

@amistre64 can you help me?

OpenStudy (amistre64):

multiply thru my 9/5 and add off the 32 and you should be good

OpenStudy (gabylovesyou):

give me the equation please so i can solve it

OpenStudy (anonymous):

\[C = \frac{5}{9} (F – 32) \\\frac{9}{5}* C = \frac{9}{5}* \frac{5}{9} (F – 32) \\ \frac{9}{5}* C =(F-32)\]

OpenStudy (anonymous):

now just add 32 to both sides

OpenStudy (anonymous):

Hope that helped.

OpenStudy (gabylovesyou):

that is for part 1?

OpenStudy (anonymous):

Yes what is the final answer.

OpenStudy (gabylovesyou):

write it how u would answer part 1 and 2 and then explain :)

OpenStudy (amistre64):

the objective of this site is not to do all the work for you, but to guide you to a solution.

OpenStudy (anonymous):

I already told you how to do it. Do you need help adding?

OpenStudy (gabylovesyou):

33.8c = f

OpenStudy (gabylovesyou):

?????

OpenStudy (anonymous):

You are way off. you really don't know how to add or multiply ?

OpenStudy (gabylovesyou):

yes but not add in that way please explain -_-

OpenStudy (anonymous):

I don't know how to teach a person how to add. I need pebbles and chalk

OpenStudy (gabylovesyou):

-_-